OpenStudy (ninabina):

The perimeter of the rectangle is 24 and the perimeter of the triangle is 23. The equations for the triangle is x+1, x+y, and y+6. The equations for the rectangle is x and 3y. Write an equations to that represents the perimeter of the triangle. Write an equation in standard form the represents the perimeter of the rectangle. Write a system of linear equations to represent the problem. The solve the system of linear equations to find x and y.

OpenStudy (anonymous):

if those equations represent the side lengths x+1+x+y+y+6 = 23 2x+2(3)y = 24 x+y=8 x+3y=12 x=8-y x=12-3y 8-y=12-3y 2y=4 y=2 x=6

OpenStudy (ninabina):

where do you get the 3 from in the first and second equations?

OpenStudy (ninabina):

never mind i see